# Last Iterate Convergence of Incremental Methods and Applications in Continual Learning

### v1

@@ 2 Incremental Gradient Descent and Shuffled SGD
We first introduce a lemma to bound the term $\sum_{t=1}^{T}\|\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\|^{2}$ in (2.1) with random permutations involved. Its proof is provided in Appendix A.
Lemma 2.6.
Under 4 and for Algorithm 1 with uniformly random (SO/RR) shuffling, we have
$\mathbb{E}\Big{[}\sum_{t=1}^{T}\|\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\|^{2}\Big{]}\leq\frac{T(T+1)}{6}\sigma_{*}^{2}.$
We then take expectation w.r.t. all the randomness on both sides of (2.1) and use Lemma 2.6. Then it remains to follow the analysis in the proof of Theorem 2.3. ∎
3 Incremental Proximal Method
In this section, we leverage the proof techniques developed in the previous section and derive the last iterate convergence bound for the incremental proximal method (IP) summarized in Algorithm 2. While the incremental proximal method is a fundamental method broadly studied in the optimization literature and thus the last iterate convergence bounds are of independent interest, our main motivation for considering this problem comes from CL applications, as discussed in the introduction. Thus we begin this section by briefly explaining this connection and reasoning.
The considered setup is motivated by the general $\ell_{2}$-regularized CL setting (Heckel, 2022; Li et al., 2023). In particular, each proximal iteration ${\bm{x}}_{k-1,t+1}=\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})$ can be interpreted as minimizing the $\ell_{2}$ (a.k.a. ridge) regularized loss $f_{t}$ corresponding to the current task $t$, which aligns with the common machine learning practice of using regularization to improve the generalization error and prevent forgetting. When $\eta_{k}\rightarrow\infty$, the proximal point step reduces to the CL setting where the learner exactly minimizes the loss of the current task, i.e., ${\bm{x}}_{k-1,t+1}=\operatorname*{arg\,min}_{{\bm{x}}\in\mathbb{R}^{d}}f_{t}({\bm{x}})$, while the regularization effect vanishes and causes larger forgetting on previous tasks. When $\eta_{k}$ is small, the proximal point step is easier to compute with larger quadratic regularization and prevents deviating from the previous iterate thus causing less forgetting. However, in this case the plasticity of the model may be deteriorated.
We also note that our analysis of the incremental proximal method is related to previous work on cyclic replays for overparameterized linear models with (S)GD (Evron et al., 2022; Swartworth et al., 2024), as (S)GD in this case acts as an implicit regularizer (Gunasekar et al., 2018; Zhang et al., 2021) (whereas proximal point update acts as an explicit regularizer). The two lines of work are not directly comparable: the prior work (Evron et al., 2022; Swartworth et al., 2024) considers exact minimization of component/task loss function and bounds the forgetting, but only addresses convex quadratics where the component loss functions have a nonempty intersecting set of minima (which implies $\sigma_{*}=0$ in 4). On the other hand, our work addresses much more general smooth convex functions (not necessarily quadratic) and does not require $\sigma_{*}=0,$ but instead relies on sufficiently large regularization.
Algorithm 2 Incremental Proximal Method (IP)
  Input: initial point ${\bm{x}}_{0}$, number of epochs $K$, step size $\{\eta_{k}\}$
  for $k=1:K$ do
     ${\bm{x}}_{k-1,1}={\bm{x}}_{k-1}$
     for $t=1:T$ do
     ${\bm{x}}_{k}={\bm{x}}_{k-1,T+1}$
  return  ${\bm{x}}_{K}$
3.1 Smooth Convex Setting
We first study the setting where the loss function of each task is convex and smooth, in contrast to prior works only focusing on nonsmooth settings (Bertsekas, 2011; Li et al., 2019, 2020). The main reasons we consider this case separately are that we can show faster convergence assuming smoothness and because it covers many regression tasks studied in prior CL works; see e.g., Evron et al. (2022, 2024). Note that with smoothness the proximal iteration is equivalent to the backward gradient step, i.e.,
${\bm{x}}_{k-1,t+1}={\bm{x}}_{k-1,t}-\eta_{k}\nabla f_{t}({\bm{x}}_{k-1,t+1}).$
Hence, much of the analysis from Section 2 can be adapted here, and the main difference lies in bounding the gap $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ within each epoch as follows.
Lemma 3.1.
Under Assumptions 1 and 3, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Algorithm 2 and for $\alpha>0,\beta>0$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the step sizes satisfy $\eta_{k}\leq\frac{1}{\sqrt{\beta}TL}$, then we have for $k\in[K]$
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\eta_{k}^{2}L\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}+\frac{\alpha}{\beta}T\big{(}f({\bm{z}})-f({\bm{x}}_{*})\big{)}$ (3.1)
@@ 3.1 Smooth Convex Setting
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{x}}_{k-1,t+1})\leq\left\langle\nabla f_{t}({\bm{x}}_{k}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\right\rangle-\frac{\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1})\|^{2}}{2L},$
$f_{t}({\bm{x}}_{k-1,t+1})-f_{t}({\bm{z}})\leq\left\langle\nabla f_{t}({\bm{x}}_{k-1,t+1}),{\bm{x}}_{k-1,t+1}-{\bm{z}}\right\rangle-\frac{\|\nabla f_{t}({\bm{x}}_{k-1,t+1})-\nabla f_{t}({\bm{z}})\|^{2}}{2L}.$
@@
Choosing ${\bm{z}}={\bm{z}}_{k-1}$ defined by (2.5) in Lemma 3.1 and handling the error terms w.r.t. $\sigma_{*}^{2}$ and $f({\bm{z}}_{k-1})-f({\bm{x}}_{*})$ as in Theorem 2.3, we can obtain the last iterate guarantee for incremental proximal methods with the same rate, as stated in the following theorem.
Theorem 3.2 (IP).
Under Assumptions 13 and 4 and for parameters $\alpha,\beta$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the constant step size $\eta\leq\frac{1}{\sqrt{\beta}TL}$, the output ${\bm{x}}_{K}$ of Algorithm 2 satisfies
@@
Follows from the proof of Theorem 2.3. ∎
A few remarks are in order here. First, the extensions to the increasing weighted averaging and RR/SO shuffling discussed in Section 2 also apply to this setting, which we omit for brevity. Second, in order to take advantage of smoothness to handle incremental updates, we require that the step size parameter $\eta$ is at most order-$1/(TL)$, which leads to less forgetting but larger regularization error. Similar step size restrictions also appear in prior work on incremental proximal method (Bertsekas, 2011; Li et al., 2019), requiring step sizes of the order $1/\sqrt{K}$ for convex Lipschitz components (same as our choice in the next subsection). On the other hand, weak regularization has only been studied asymptotically in Evron et al. (2023) for linear classification tasks with the regularization parameter tending to zero, in the need of relating to the sequential max-margin projections for their analysis. Evron et al. (2023) also left a concluding remark that another challenging but rewarding avenue is to extend our analysis to finite regularization strengths. Hence, at the current stage, it is unclear whether such a restriction on our small step size can be lifted. An interesting aspect of our result is that there is a critical value
$\eta_{*}=\min\Big{\{}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2/3}}{2^{1/3}T\sigma_{*}^{2/3}L^{1/3}K^{1/3}(1+\beta/\alpha)^{1/3}},\,\frac{1}{\sqrt{\beta}TL}\Big{\}}$
such that if decrease $\eta$ beyond $\eta_{*}$, both the regularization error and the forgetting increase. It is thus an interesting open question to characterize how the forgetting is affected for $\eta<\eta_{*}$ (i.e., when we decrease the regularization error). Lastly, we note that for $\ell_{2}$-regularized CL, the role of the regularization parameter in the trade-off between the forgetting and the regularization error has only been theoretically calibrated for the setting of two linear regression tasks without cyclic replay (Heckel, 2022; Li et al., 2023), while the extensions to the cyclic replays or the general convex tasks remain unknown.
3.2 Convex Lipschitz Setting
We now further relax the smoothness assumption and consider the convex Lipschitz setting, with applications such as linear classification tasks considered in Evron et al. (2023).
To carry out the analysis, we leverage the standard fact that the proximal iteration is equivalent to the gradient step w.r.t. the Moreau envelope, i.e.,
while the gradient of the Moreau envelope $\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})$ belongs to the subdifferential $\partial f_{t}({\bm{x}}_{k-1,t+1})$, thus is bounded by the Lipschitz constant of $f_{t}$. We use these observations to derive an upper bound on $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ for each epoch, as summarized in the following lemma.
Lemma 3.3.
Under Assumptions 1 and 2, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Algorithm 2, we have for $k\in[K]$
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}+\frac{T(T-1)G^{2}\eta_{k}}{2}.$ (3.2)

### v2

@@ Appendix A Omitted Proofs From Section 2
We first introduce a lemma to bound the term $\sum_{t=1}^{T}\|\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\|^{2}$ in Lemma 2.1 with random permutations involved. Its proof is provided in Appendix A.
Lemma A.1.
Under Assumptions 4 and for Alg. 1 with uniformly random (SO/RR) shuffling, we have
$\textstyle\mathbb{E}\big{[}\sum_{t=1}^{T}\|\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\|^{2}\big{]}\leq\frac{T(T+1)}{6}\sigma_{*}^{2}.$
We then take expectation w.r.t. all the randomness on both sides of the inequality in Lemma 2.1 and use Lemma A.1. Then it remains to follow the analysis in the proof of Theorem 2.3. ∎
See A.1
Proof.
=\;= $\frac{T(T+1)}{6}\sigma_{*}^{2},$
Appendix B Omitted Proofs from Section 3
B.1 Convex Smooth Setting
Lemma B.1.
Under Assumptions 1 and 3, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Alg. 2 and for $\alpha>0,\beta>0$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the step sizes satisfy $\eta_{k}\leq\frac{1}{\sqrt{\beta}TL}$, then we have for $k\in[K]$
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq$ $\eta_{k}^{2}L\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}+\frac{\alpha}{\beta}T\big{(}f({\bm{z}})-f({\bm{x}}_{*})\big{)}$ (B.1)
@@ B.1 Convex Smooth Setting
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{x}}_{k-1,t+1})\leq\left\langle\nabla f_{t}({\bm{x}}_{k}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\right\rangle-\frac{1}{2L}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1})\|^{2},$
$f_{t}({\bm{x}}_{k-1,t+1})-f_{t}({\bm{z}})\leq\left\langle\nabla f_{t}({\bm{x}}_{k-1,t+1}),{\bm{x}}_{k-1,t+1}-{\bm{z}}\right\rangle-\frac{1}{2L}\|\nabla f_{t}({\bm{x}}_{k-1,t+1})-\nabla f_{t}({\bm{z}})\|^{2}.$
@@
See 3.2
@@
For all parts of the proof, we consider $1$-dimensional quadratics
$f(x):=\frac{1}{T}\sum_{t=1}^{T}f_{t}(x),\quad\text{where}\quad f_{t}(x)=\frac{L}{2}(x-\delta_{t})^{2}$
for $t\in[T]$, $L>0$, and appropriately chosen sequences of $\{\delta_{t}\}_{t\in[T]}\subseteq\mathbb{R}$.
It is immediate that $f(x)$ is minimized at $x_{*}=\frac{1}{T}\sum_{t=1}^{T}\delta_{t}$. Observe that Alg. 2 using a constant step size $\eta>0$ has closed-form updates on $f$, i.e.,
$x_{k+1}=\gamma^{n}x_{k}+(1-\gamma)\sum_{t=1}^{T}\gamma^{T-t}\delta_{t},$
where $\gamma=\frac{1}{\eta L+1}\in(0,1)$. Given any initial point $x_{0}$, by iterating we have
$\textstyle x_{k}-x_{*}$ $=\gamma^{kT}x_{0}+\sum_{t=1}^{T}\Big{(}\frac{\gamma^{T-t}(1-\gamma)(1-\gamma^{kT})}{1-\gamma^{T}}-\frac{1}{T}\Big{)}\delta_{t}$ (B.2)
$\overset{k\rightarrow\infty}{\longrightarrow}\sum_{t=1}^{T}\Big{(}\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\Big{)}\delta_{t}.$ (B.3)
  1. 1.
    Consider the weight $\delta_{T}$ in Eq. (B.3). Since $\gamma\in(0,1)$, we have
    $\frac{1-\gamma}{1-\gamma^{T}}-\frac{1}{T}=\frac{1}{\sum_{t=0}^{T-1}\gamma^{t}}-\frac{1}{T}>0.$
    Then for any $\{\delta_{t}\}_{t\in[T]}$ such that $\mathrm{sgn}(\delta_{t})=\mathrm{sgn}\big{(}\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\big{)}$ and $\delta_{T}>0$, we know that
    $\lim_{k\rightarrow\infty}f(x_{k})-f(x_{*})\overset{(i)}{=}\frac{L}{2}\lim_{k\rightarrow\infty}(x_{k}-x_{*})^{2}\geq\frac{L}{2}\Big{(}\frac{1-\gamma}{1-\gamma^{T}}-\frac{1}{T}\Big{)}^{2}\delta_{T}^{2}>0,$
    where $(i)$ is due to $f$ being both $L$-strongly convex and $L$-smooth.
  2. 2.
    Consider the weights of $\delta_{t}$ in Eq. (B.3). Since $\gamma\in(0,1)$, we have
    $0\leq\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}\leq\gamma^{T-t}\leq 1,$
    thus for any $t\in[T-1]$
    $\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\in\Big{[}-\frac{1}{T},\frac{T-1}{T}\Big{)}.$
    For $t=T$, given any fixed $\gamma<1$, we have
    $\frac{1-\gamma}{1-\gamma^{T}}-\frac{1}{T}=\frac{1}{\sum_{t=0}^{T-1}\gamma^{t}}-\frac{1}{T}>0.$
    Hence, for the sequence $\{\delta_{t}\}$ such that
    $|\delta_{t}|<\frac{T\sqrt{2/L}}{(T-1)^{2}}\;(t\in[T-1]),\quad\delta_{T}>\frac{2\sqrt{2/L}}{\frac{1-\gamma}{1-\gamma^{T}}-\frac{1}{T}},$
    then combining the bounds on the weights of $\delta_{t}$ with Eq. (B.3) we obtain
    $\lim_{k\rightarrow\infty}x_{k}-x_{*}=\;$ $\sum_{t=1}^{T}\Big{(}\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\Big{)}\delta_{t}$
[…]
    >\;> $\sqrt{2/L}.$
    Since $f$ is $L$-smooth and $L$-strongly convex, we know that
    $\lim_{k\rightarrow\infty}f(x_{k})-f(x_{*})=\frac{L}{2}\lim_{k\rightarrow\infty}(x_{k}-x_{*})^{2}>1,$
    thus finishing the proof of the second part. We note in passing that 1 on the right-hand side can be replaced by any constant using a simple rescaling.
  3. 3.
    Observe that given a fixed step size $\eta>0$, we can choose a sequence $\{\delta_{t}\}_{t\in[T]}$ such that $\mathrm{sgn}(\delta_{t})=\mathrm{sgn}\big{(}\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\big{)}$ for all $t\in[T]$, thus for any initial point $x_{0}\geq x_{*}$:
    $f(x_{k})-f(x_{*})=\;$ $\frac{L}{2}\Big{(}\gamma^{kT}(x_{0}-x_{*})+(1-\gamma^{kT})\sum_{t=1}^{T}\Big{(}\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\Big{)}\delta_{t}\Big{)}^{2}$
[…]
    $+(1-\gamma^{kT})^{2}L\sum_{s\neq t\in[T]}\Big{(}\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\Big{)}\Big{(}\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\Big{)}\delta_{s}\delta_{t}.$
    Without loss of generality, taking a sufficiently large $k\geq\frac{\log_{\gamma}(1-2/\sqrt{5})}{T}$, we obtain
    $f(x_{k})-f(x_{*})\geq\frac{2L}{5}\sum_{t=1}^{T}\Big{(}\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\Big{)}^{2}\delta_{t}^{2}.$
    Then for any step size $\eta\geq 1/TL$, we have $\gamma=\frac{1}{\eta L+1}\leq\frac{T}{T+1}$. Consider $t=T$, we can bound
    $\frac{1-\gamma}{1-\gamma^{T}}-\frac{1}{T}\geq\frac{1-\frac{T}{T+1}}{1-(\frac{T}{T+1})^{T}}-\frac{1}{T}=\frac{1}{T}\big{(}\frac{1}{1-(1-\frac{1}{T})^{T}}-1\big{)}>\frac{\mathrm{e}-1}{T}.$
    Thus, for $\delta_{T}\geq\frac{\sqrt{5}T\sqrt{\varepsilon}}{\sqrt{2}(\mathrm{e}-1)L}$, we have
    $f(x_{k})-f(x_{*})\geq\frac{2L}{5}\Big{(}\frac{1-\gamma}{1-\gamma^{T}}-\frac{1}{T}\Big{)}^{2}\delta_{T}^{2}>\varepsilon.$
    On the other hand, recalling the definition in Assumption 4, we have in this example that
    $\sigma_{*}^{2}=\frac{L^{2}}{T}\sum_{t=1}^{T}\Big{(}\frac{\sum_{t=1}^{T}\delta_{t}}{T}-\delta_{t}\Big{)}^{2}=\frac{L^{2}}{T}\Big{(}\frac{(\sum_{t=1}^{T}\delta_{t})^{2}}{T}-2\frac{(\sum_{t=1}^{T}\delta_{t})^{2}}{T}+\sum_{t=1}^{T}\delta_{t}^{2}\Big{)}\leq L^{2}\sum_{t=1}^{T}\delta_{t}^{2}/T.$
    Thus for any step size $\eta\geq\frac{16\sqrt{\varepsilon}}{\sqrt{TL}\sigma_{*}}\geq\frac{16\sqrt{\varepsilon}}{L^{3/2}\sqrt{\sum_{t=1}^{T}\delta_{t}^{2}}}$, we have $\gamma=\frac{1}{\eta L+1}\leq\frac{\sqrt{\sum_{t=1}^{T}\delta_{t}^{2}}}{16\sqrt{\varepsilon/L}+\sqrt{\sum_{t=1}^{T}\delta_{t}^{2}}}$.
    We now proceed by bounding the weight of $\delta_{T}$. In particular, let $\gamma=1-\kappa$ for some $\kappa>0$, and assume that $\kappa\leq\frac{1}{T+1}<\frac{1}{T}$ without loss of generality by the discussion above. Since $\kappa T<1$ and $T\geq 2$, we have
    $(1-\kappa)^{T}\geq 1-\kappa T+\frac{\kappa^{2}T(T-1)}{4},$
    which leads to
    $1-\gamma^{T}=1-(1-\kappa)^{T}\leq\kappa T-\frac{\kappa^{2}T(T-1)}{4}.$
    Hence, we have
    $\frac{1-\gamma}{1-\gamma^{T}}\geq\frac{\kappa}{\kappa T-\frac{\kappa^{2}T(T-1)}{4}}=\frac{1}{T}\frac{1}{1-\kappa(T-1)/4}.$
    Further noticing that
    $\frac{1}{1-\kappa(T-1)/4}\geq 1+\frac{\kappa(T-1)}{4}\geq 1+\frac{\kappa T}{8}$
    for $T\geq 2$ and $\kappa<\frac{1}{T}$, then we have
    $\frac{1-\gamma}{1-\gamma^{T}}-\frac{1}{T}\geq\frac{1}{T}(1+\frac{\kappa T}{8})-\frac{1}{T}=\frac{\kappa}{8}.$
    Recall that $\gamma\leq\frac{\sqrt{\sum_{t=1}^{T}\delta_{t}^{2}}}{16\sqrt{\varepsilon/L}+\sqrt{\sum_{t=1}^{T}\delta_{t}^{2}}}$, then we obtain
    $\frac{1-\gamma}{1-\gamma^{T}}-\frac{1}{T}\geq\frac{2\sqrt{\varepsilon/L}}{16\sqrt{\varepsilon/L}+\sqrt{\sum_{t=1}^{T}\delta_{t}^{2}}}\geq\frac{\sqrt{3\varepsilon/L}}{\sqrt{\sum_{t=1}^{T}\delta_{t}^{2}}},$
    for $\{\delta_{t}\}_{t\in[T]}$ such that $\sqrt{\sum_{t=1}^{T}\delta_{t}^{2}}\geq 16\sqrt{3}(2+\sqrt{3})\sqrt{\varepsilon/L}$. Thus, for the sequence $\{\delta_{t}\}_{t\in[T]}$ such that $\delta_{T}^{2}>\frac{5}{6}\sum_{t=1}^{T}\delta_{t}^{2}$ and $\mathrm{sgn}(\delta_{t})=\mathrm{sgn}\big{(}\frac{\gamma^{T-t}(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\big{)}$, we have
    $f(x_{k})-f(x_{*})\geq\;$ $\frac{2L}{5}\Big{(}\frac{(1-\gamma)}{1-\gamma^{T}}-\frac{1}{T}\Big{)}^{2}\delta_{T}^{2}$
    >\;> $\frac{2L}{5}\frac{3\varepsilon\delta_{T}^{2}}{L\sum_{t=1}^{T}\delta_{t}^{2}}$
    >\;> $\varepsilon,$
completing the proof. ∎
B.2 Convex Lipschitz Setting
Lemma B.2.
Under Assumptions 1 and 2, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Alg. 2, we have for $k\in[K]$
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq$ $\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}+\frac{T(T-1)G^{2}\eta_{k}}{2}.$ (B.4)